1.algebra the measures of two supplementary angles are 4r+7 and r-2, find the measures of the angles.
find the distance between each pair of points.
1.G(0,0), H(-3,4) 2.N(5,2), K(-2,8) 3.A(-4,-4), W(-2,2)
1
Expert's answer
2011-10-14T12:31:28-0400
A=4r+7, B=r-2

A+B=5r+5=180->5r=175->r=35. So A=147, B=33.

GH=HG=sqrt((-3)^2+4^2)=5

NK=KN=sqrt((-2-5)^2+(8-2)^2)=9.22

AW=WA=sqrt((-2-(-4))^2+(2-(-4))^2)=6.324